# Right on Time: Revising Time Series Models by Constraining their Explanations

### v2

@@ 5 Conclusion
In this work, we present Right on Time a method to mitigate confounding factors in time series data with the help of human feedback. By revising the model, RioT significantly diminishes the influence of these factors, steering the model to align with the correct reasons. Using popular time series models on several manually confounded datasets and the newly introduced, naturally confounded, real-world dataset P2S showcases that they are indeed subject to confounders. Our results, however, demonstrate that applying RioT to these models can mitigate confounders in the data. Furthermore, we have unveiled that addressing solely the time domain is insufficient for revising the model to focus on the correct reasons, which is why we extended our method beyond it. Feedback in the frequency domain provides an additional way to steer the model away from confounding factors and towards the right reasons.
Extending the application of RioT to multivariate time series represents a logical next step, and exploring the integration of various explainer types is another promising direction. Additionally, we aim to apply RioT, especially RioTfreq, to other modalities as well, offering a more nuanced approach to confounder mitigation. It should be noted that while our method shows potential in its current iteration, interpreting attributions in time series data remains a general challenge.
Impact Statement
@@
Acknowledgment
This work received funding by the EU project EXPLAIN, funded by the Federal Ministry of Education and Research (grant 01—S22030D). Additionally, this was funded by the project ”The Adaptive Mind” from the Hessian Ministry of Science and the Arts (HMWK), the ”ML2MT” project from the Volkswagen Stiftung, and the Priority Program (SPP) 2422 in the subproject “Optimization of active surface design of high-speed progressive tools using machine and deep learning algorithms“ funded by the German Research Foundation (DFG). The latter also contributed the data for the P2S dataset. Furthermore, this work benefited from the HMWK project ”The Third Wave of Artificial Intelligence - 3AI”.
Appendix A Appendix
A.1 Implementation Details
To compute explanations with Integrated Gradients, we follow the common practice of using a baseline of zeros. The standard approach worked well in our experiments, thus we did not explore other baseline choices in this work.
To find suitable parameters for model training, we performed a hyperparameter search over batch size, learning rate and number of training epochs. We then used these parameters for all model trainings and evaluations, with and without RioT. We furthermore selected suitable $\lambda$ values for RioT with a hyperparameter selection on the respective validation sets. The exact values for the model training parameters and the $\lambda$ values can be found in the provided code.
@@ A.1 Implementation Details
A.2 UCR Dataset selection
@@
We focused our evaluation on a subset of UCR datasets with a minimum size. Our selection process was as follows: First, we discarded all multivariate datasets, as we only considered univariate data in this paper. Then we removed all datasets of the category SIMULATED, to avoid datasets which were synthetic from the beginning. We furthermore considered only datasets with less than 10 classes, as having a per-class confounder on more than 10 classes would result in a very high number of different confounders, which would probably rarely happen. Besides these criteria, we discarded all datasets with less than 1000 training samples or a per sample length of less than 100, to avoid the small datasets of UCR. Lastly, we discarded all datasets which could not be loaded, which lead to the resulting four datasets: Fault Detection A, Ford A, Ford B and Sleep.
A.3 Additional Experimental Results
@@
Energy (Forecasting MAE $\downarrow$) Train Test
TiDEReference 0.40 ±0.01 0.38 ±0.02
@@
TiDEfreq,sp + RioTfreq,sp 0.39 ±0.01 0.55 ±0.01
This section of the appendix provides further insights into our forecasting experiments, specifically showcasing the performance through the MAE metric. The accompanying Tab. 5 details the MAE outcomes for various configurations of our model, emphasizing the effectiveness of integrating spatial and frequency feedback via RioT. These results complement our primary findings (cf. Tab. 4), illustrating the model’s enhanced forecasting performance when addressing confounders across both the time and frequency domains.
A.4 Confounded Dataset from a High-speed Progressive Tool
@@
A.4.1 Real-World setting
@@
A.4.2 Data collection
An indirect piezoelectric force sensor (Kistler 9240A) was integrated into the upper mould mounting plate of the deep-drawing stage for data acquisition. The sensor is located directly above the punch and records not only the indirect process force but also the blank holder forces which are applied by spring assemblies between the upper mounting plate and the blank holder plate. The data is recorded at a sampling frequency of 25 kHz. The material used is DC04 with a width of 50 mm and a thickness of 0.5 mm. The voltage signals from the sensors are digitised using a ”CompactRIO” (NI cRIO 9047) with integrated NI 9215 measuring module (analogue voltage input $\pm$ 10 V). Data recording is started via an inductive proximity switch when the press ram passes below a defined stroke height during the stroke movement and is stopped again as it passes the inductive proximity switch during the return stroke movement. Due to the varying process speed caused by the stroke speeds that have been set, the recorded time series have a different number of data points. Further, there are slight variations in the length of the time series withing one experiment. For this reason, all time series are interpolated to a length of 4096 data points and we discard any time series that deviate considerably from the mean length of a run (i.e., threshold of 3). A total of 12 series of experiments, shown in Tab. 6, were carried out with production rates from 80 to 225 spm. To simulate a defect, the spring hardness of the blank holder was manipulated in the test series that were marked as defect. The manipulated experiments result in the component bursting and tearing during production. In a real production environment, this would lead directly to the parts being rejected.
A.4.3 Data characteristics
@@ A.4.4 Confounders
In the experiment setting of Tab. 3, the training data set is selected in such a way that the stroke rate correlates with the class label, i.e., there are only normal experiments with small stroke rates and defect ones with high stroke rate. Experiment 1, 2, 3, 10, 11, 12 are the training data and the remaining experiments are the test data. To get a reference setting where the model is not affected by any confounder, we use normal and defect experiments with the same speed in training and respectively test data. This results in experiments 1, 3, 5, 7, 9, 11 in the training set and the remaining in the test set.

### v3

@@ 5 Conclusion
In this work, we present Right on Time a method to mitigate confounding factors in time series data with the help of human feedback. By revising the model, RioT significantly diminishes the influence of these factors, steering the model to align with the correct reasons. Using popular time series models on several manually confounded datasets and the newly introduced, naturally confounded, real-world dataset P2S showcases that they are indeed subject to confounders. Our results, however, demonstrate that applying RioT to these models can mitigate confounders in the data. Furthermore, we have unveiled that addressing solely the time domain is insufficient for revising the model to focus on the correct reasons, which is why we extended our method beyond it. Feedback in the frequency domain provides an additional way to steer the model away from confounding factors and towards the right reasons. Extending the application of RioT to multivariate time series represents a logical next step, and exploring the integration of various explainer types is another promising direction. Additionally, we aim to apply RioT, especially RioTfreq, to other modalities as well, offering a more nuanced approach to confounder mitigation. It should be noted that while our method shows potential in its current iteration, interpreting attributions in time series data remains a general challenge.
Acknowledgment
This work received funding by the EU project EXPLAIN, funded by the Federal Ministry of Education and Research (grant 01—S22030D). Additionally, it was funded by the project "The Adaptive Mind" from the Hessian Ministry of Science and the Arts (HMWK), the "ML2MT" project from the Volkswagen Stiftung, and the Priority Program (SPP) 2422 in the subproject “Optimization of active surface design of high-speed progressive tools using machine and deep learning algorithms“ funded by the German Research Foundation (DFG). The latter also contributed the data for P2S. Furthermore, this work benefited from the HMWK project ”The Third Wave of Artificial Intelligence - 3AI”.
Appendix A Appendix
A.1 Impact Statement
@@
A.2 Implementation and Experimental Details
Adaption of Integrated Gradients (IG). A part of IG is a multiplication of the model gradient with the input itself, improving the explanation’s quality [Shrikumar et al., 2017]. However, this multiplication makes some implicit assumptions about the input format. In particular, it assumes that there are no inputs with negative values. Otherwise, multiplying the attribution score with a negative input would flip the attribution’s sign, which is not desired. For images, this is unproblematic because they are always equal to or larger than zero. In time series, negative values can occur and normalization to make them all positive is not always suitable. To avoid this problem, we use only the input magnitude and not the input sign to compute the IG attributions.
Computing Explanations. To compute explanations with Integrated Gradients, we followed the common practice of using a baseline of zeros. The standard approach worked well in our experiments, so we did not explore other baseline choices in this work. For the implementation, we utilized the widely-used Captum [^3] https://github.com/pytorch/captum library, where we patched the captum._utils.gradient.compute_gradients function to allow for the propagation of the gradient with respect to the input to be propagated back into the parameters.
Model Training and Hyperparameters. To find suitable parameters for model training, we performed a hyperparameter search over batch size, learning rate, and the number of training epochs. We then used these parameters for all model trainings and evaluations, with and without RioT. In addition, we selected suitable $\lambda$ values for RioT with a hyperparameter selection on the respective validation sets. The exact values for the model training parameters and the $\lambda$ values can be found in the provided code.
@@ A.2 Implementation and Experimental Details
All experiments were executed using our Python 3.11 and PyTorch code, which is available in the provided code. To ensure reproducibility and consistency, we utilized Docker. Configurations and Python executables for all experiments are provided in the repository.
Hardware. To conduct our experiments, we utilized single GPUs from Nvidia DGX2 machines equipped with A100-40G and A100-80G graphics processing units.
By maintaining a consistent hardware setup and a controlled software environment, we aimed to ensure the reliability and reproducibility of our experimental results.
A.3 UCR Dataset selection
@@
We focused our evaluation on a subset of UCR datasets with a minimum size. Our selection process was as follows: First, we discarded all multivariate datasets, as we only considered univariate data in this paper. Then we removed all datasets with time series of different length or missing values. We further excluded all datasets of the category SIMULATED, to avoid datasets which were synthetic from the beginning. We furthermore considered only datasets with less than 10 classes, as having a per-class confounder on more than 10 classes would result in a very high number of different confounders, which would probably rarely happen. Besides these criteria, we discarded all datasets with less than 1000 training samples or a per sample length of less than 100, to avoid the small datasets of UCR, which leads to the resulting four datasets: Fault Detection A, Ford A, Ford B and Sleep.
A.4 Computational Costs of RioT
Training a model with RioT induces additional computational costs. The right-reason term requires computations of additional gradients. Given a model $f_{\theta}(x)$, parameterized by $\theta$ and input $x$, then computing the right reason loss with a gradient-based explanation method requires the computation of the mixed-partial derivative $\frac{\partial^{2}f_{\theta}(x)}{\partial\theta\partial x}$, as a gradient-based explanation includes the derivative $\frac{\partial f_{\theta}(x)}{\partial x}$. While this mixed partial derivative is a second order derivative, this does not substantially increase the computational costs of our method for two main reasons. First, we are never explicitly materializing the Hessian matrix. Second, the second-order component of our loss can be formulated as a Hessian-vector product:
$\frac{\partial\mathcal{L}}{\partial\theta}=g+\frac{\lambda}{2}H_{\theta x}(e(x)-a(x))$ (6)
where $g=\frac{\partial\mathcal{L}_{\mathrm{RA}}}{\partial\theta}$ is the partial derivative of the right answer loss and if $H$ is the full joint Hessian matrix of the loss with respect to $\theta$ and $x$, then $H_{\theta x}$ is the sub-block of this matrix mapping $x$ into $\theta$ (cf. Fig. 6), given by $H_{\theta x}=\frac{\partial^{2}f_{\theta}(x)}{\partial\theta\partial x}$. Hessian-vector products are known to be fast to compute [Martens, 2010], enabling the right-reason loss computation to scale to large models and inputs.
[…]
Figure 6: Illustration of the Hessian matrix with its respective sub-blocks. The mapping from $x$ into $\theta$ is highlighted in blue.
A.5 Details on Confounding Factors
In the datasets which are not P2S, we added synthetic confounders to evaluate the effectiveness of confounders. In the following, we provide details on the nature of these confounders in the four settings:
Classification Spatial. For classification datasets, spatial confounders are specific patterns for each class. The pattern is added to every sample of that class in the training data, resulting in a spurious correlation between the pattern and the class label. Specifically, we replace $T$ time steps with a sine wave according to:
[…]
while $t\in\{0,1,\dots,T\}$ and $j$ represents the class index, simulating a spurious correlation between the confounder and class index.
Classification Frequency. Similar to the spatial case, frequency confounders for classification are specific patterns added to the entire series, altering all time steps by a small amount. The confounder is represented as a sine wave and is applied additively to the full sequence ($T=S$):
[…]
where $A$ resembles the confounder amplitude.
Forecasting Spatial. For forecasting datasets, spatial confounders are shortcuts that act as the actual solution to the forecasting problem. Periodically, data from the time series is copied back in time. This “back-copy” is a shortcut for the forecast, as it resembles the time steps of the forecasting window. Due to the windowed sampling from the time series, this shortcut occurs at every second sample. The exact confounder formulation is outlined in the sketch below (Fig. 7), with an exemplary lookback length of $9$, forecasting horizon of $3$ and window stride of $6$. This results in a shortcut confounder in samples 1 and 3 (marked red) and overlapping in sample 2 (marked orange).
1. SampleLookbackHorizonUnconfounded:01234567891011Confounded:9101134567891011Feedback:1110000002. SampleUnconfounded:67891011121314151617Confounded:67891011212223151617Feedback:0000000003. SampleUnconfounded:121314151617181920212223Confounded:212223151617181920212223Feedback:111000000Confounder:Overlapping Confounder:
Figure 7: Schematic overview of how the time series were confounded in the spatial forecasting experiments
Forecasting Frequency. This setting differs from the previous shortcut confounders. The frequency confounder for forecasting is a recurring Dirac impulse with a certain frequency, added every $k$ time steps over the entire sequence (of length $S$), including the forecasting windows. This impulse is present throughout all of the training data, distracting the model from the real forecast. The confounder is present at all time steps: $i\in\{n\cdot k|n\in\mathbb{N}\wedge n\cdot k\leq S\}$ with a strength of $A$:
[…]
In conclusion, confounders are only present in the training data, not validation or test data. We generate an annotation mask based on the confounder area or frequency to simulate human feedback. This mask is applied to all confounded samples except in our feedback scaling experiment.
A.6 Additional Experimental Results
Table 5: Feedback percentage for forecasting across all datasets, reported for the TiDE model. Corresponding to (test) results shown in Fig. 5, a higher percentage indicates more feedback, lower is better.
Metric Feedback ETTM1 Energy Weather
Spatial Freq Spatial Freq Spatial Freq
MAE ($\downarrow$) 0% 0.54 ±0.01 0.74 ±0.06 0.85 ±0.01 0.53 ±0.07 0.29 ±0.01 0.49 ±0.09
5% 0.52 ±0.00 0.63 ±0.03 0.62 ±0.01 0.40 ± 0.02 0.28 ±0.01 0.43 ±0.03
10% 0.52 ±0.00 0.63 ±0.03 0.61 ±0.01 0.40 ± 0.02 0.27 ±0.01 0.43 ±0.03
25% 0.52 ±0.00 0.63 ±0.03 0.58 ±0.01 0.41 ±0.01 0.25 ±0.01 0.43 ±0.04
50% 0.52 ±0.00 0.63 ±0.03 0.57 ± 0.01 0.41 ±0.01 0.24 ± 0.01 0.44 ±0.05
75% 0.52 ±0.01 0.63 ±0.03 0.57 ± 0.01 0.41 ±0.01 0.24 ± 0.01 0.45 ±0.06
100% 0.51 ± 0.01 0.60 ± 0.05 0.58 ±0.01 0.40 ± 0.03 0.24 ± 0.01 0.41 ± 0.02
MSE ($\downarrow$) 0% 0.54 ±0.03 0.69 ±0.08 1.19 ±0.03 0.34 ±0.08 0.15 ±0.01 0.31 ±0.09
5% 0.54 ±0.01 0.52 ±0.03 0.60 ±0.02 0.20 ± 0.01 0.14 ±0.01 0.24 ±0.02
10% 0.53 ±0.01 0.52 ±0.03 0.57 ±0.02 0.20 ± 0.01 0.14 ±0.01 0.24 ±0.02
25% 0.53 ±0.01 0.52 ±0.03 0.53 ±0.02 0.22 ±0.01 0.11 ± 0.01 0.24 ±0.03
50% 0.53 ±0.01 0.52 ±0.03 0.51 ± 0.02 0.22 ±0.01 0.11 ± 0.01 0.25 ±0.04
75% 0.52 ±0.01 0.51 ±0.03 0.52 ±0.02 0.22 ±0.01 0.11 ± 0.01 0.26 ±0.05
100% 0.48 ± 0.01 0.49 ± 0.07 0.52 ±0.02 0.21 ±0.02 0.11 ± 0.01 0.22 ± 0.01
Table 6: Feedback percentage for classification across all datasets, reported for the FCN model. Corresponding to results shown in Fig. 5, a higher percentage indicates more feedback, higher is better.
Feedback Fault Detection A (ACC $\uparrow$) FordA (ACC $\uparrow$) FordB (ACC $\uparrow$) Sleep (ACC $\uparrow$)
Spatial Freq Spatial Freq Spatial Freq Spatial Freq
0% 0.74 ±0.06 0.87 ±0.03 0.71 ±0.08 0.73 ±0.01 0.63 ±0.03 0.60 ±0.01 0.10 ±0.03 0.27 ±0.02
5% 0.88 ±0.00 0.88 ±0.01 0.81 ±0.03 0.80 ±0.03 0.66 ±0.03 0.66 ± 0.02 0.53 ±0.03 0.49 ± 0.00
10% 0.89 ±0.02 0.89 ±0.01 0.82 ±0.04 0.79 ±0.02 0.66 ±0.03 0.64 ±0.03 0.48 ±0.09 0.48 ±0.02
25% 0.92 ±0.01 0.89 ±0.01 0.83 ±0.02 0.78 ±0.01 0.67 ±0.02 0.65 ±0.01 0.49 ±0.08 0.42 ±0.08
50% 0.95 ± 0.01 0.88 ±0.01 0.82 ±0.03 0.81 ±0.05 0.67 ±0.02 0.65 ±0.00 0.55 ± 0.03 0.44 ±0.07
75% 0.95 ± 0.01 0.88 ±0.01 0.81 ±0.03 0.80 ±0.04 0.65 ±0.03 0.64 ±0.00 0.54 ±0.04 0.44 ±0.07
100% 0.93 ±0.03 0.90 ± 0.03 0.84 ± 0.02 0.83 ± 0.02 0.68 ± 0.02 0.65 ±0.01 0.54 ±0.05 0.45 ±0.07
Table 7: RioT can successfully overcome confounders in time series forecasting. MAE values on the confounded training set and the unconfounded test set with Unconfounded being the ideal scenario where the model is not affected by any confounder.
Model Config (MAE $\downarrow$) ETTM1 Energy Weather
Train Test Train Test Train Test
NBEATS Unconfounded 0.39 ±0.01 0.48 ±0.01 0.44 ±0.02 0.38 ±0.01 0.21 ±0.01 0.12 ±0.01
SP Conf 0.34 ±0.01 0.54 ±0.01 0.44 ±0.03 0.77 ±0.01 0.21 ±0.01 0.30 ±0.04
+ RioTsp 0.40 ±0.01 $\bullet$ 0.52 ±0.01 0.53 ±0.02 $\bullet$ 0.62 ±0.01 0.23 ±0.01 $\bullet$ 0.22 ±0.01
Freq Conf 0.39 ±0.01 0.47 ±0.01 0.45 ±0.03 0.45 ±0.03 0.21 ±0.03 0.45 ±0.06
+ RioTfreq 0.40 ±0.01 $\bullet$ 0.47 ±0.01 0.45 ±0.03 $\bullet$ 0.44 ±0.02 0.59 ±0.22 $\bullet$ 0.39 ±0.01
PatchTST Unconfounded 0.50 ±0.01 0.49 ±0.01 0.39 ±0.00 0.38 ±0.01 0.38 ±0.03 0.18 ±0.00
SP Conf 0.46 ±0.00 0.53 ±0.01 0.41 ±0.00 0.78 ±0.01 0.32 ±0.04 0.33 ±0.00
+ RioTsp 0.46 ±0.01 $\bullet$ 0.52 ±0.01 0.51 ±0.00 $\bullet$ 0.53 ±0.01 0.54 ±0.12 $\bullet$ 0.28 ±0.00
Freq Conf 0.53 ±0.01 0.81 ±0.07 0.15 ±0.00 0.64 ±0.03 0.58 ±0.03 0.41 ±0.05
+ RioTfreq 0.92 ±0.05 $\bullet$ 0.80 ±0.02 0.97 ±0.86 $\bullet$ 0.57 ±0.02 0.65 ±0.01 $\bullet$ 0.40 ±0.01
TiDE Unconfounded 0.36 ±0.01 0.48 ±0.01 0.40 ±0.01 0.38 ±0.02 0.36 ±0.02 0.13 ±0.00
SP Conf 0.32 ±0.01 0.54 ±0.01 0.40 ±0.01 0.85 ±0.01 0.32 ±0.03 0.29 ±0.01
+ RioTsp 0.34 ±0.01 $\bullet$ 0.51 ±0.01 0.57 ±0.01 $\bullet$ 0.58 ±0.01 0.35 ±0.03 $\bullet$ 0.24 ±0.01
Freq Conf 0.18 ±0.01 0.74 ±0.06 0.18 ±0.01 0.53 ±0.07 0.65 ±0.05 0.49 ±0.09
+ RioTfreq 0.19 ±0.01 $\bullet$ 0.60 ±0.05 0.18 ±0.01 $\bullet$ 0.40 ±0.03 0.79 ±0.16 $\bullet$ 0.41 ±0.02
@@ A.6 Additional Experimental Results
Energy (MAE $\downarrow$) Train    Test
TiDEUnconfounded 0.40 ±0.01 0.38 ±0.02
@@
TiDEfreq,sp + RioTfreq,sp 0.39 ±0.01 $\bullet$ 0.55 ±0.01
This section provides further insights into our experiments, covering both forecasting and classification tasks. Specifically, it showcases performance through various metrics such as MAE, MSE, and accuracy, and explores different feedback configurations.
Feedback Generalization.: Tab. 6 and Tab. 5 detail provided feedback percentages for forecasting and classification across all datasets, respectively. These tables report the performance of the TIDE and FCN models, highlighting how different levels of feedback impact model outcomes on various datasets. Tab. 5 focuses on MAE and MSE for forecasting, while Tab. 6 presents ACC for classification.
Removing Confounders for Time Series Forecasting. Tab. 7 reports the MAE results for our forecasting experiment across different models, datasets and configurations. It emphasizes how well each model performs on both the confounded training set and after applying RioT, with the Unconfounded configuration representing the ideal scenario unaffected by confounders.
Removing Multiple Confounders at Once. Tab. 8 reports the MAE values and illustrates the effectiveness of combining spatial and frequency feedback via RioT for the TiDE model. The results demonstrate significant improvements in forecasting accuracy when integrating both feedback domains compared to using them separately.
Appendix B Confounded Dataset from a High-speed Progressive Tool
@@
B.1 Real-World setting
@@
B.2 Data collection
An indirect piezoelectric force sensor (Kistler 9240A) was integrated into the upper mould mounting plate of the deep-drawing stage for data acquisition. The sensor is located directly above the punch and records not only the indirect process force but also the blank holder forces which are applied by spring assemblies between the upper mounting plate and the blank holder plate. The data is recorded at a sampling frequency of 25 kHz. The material used is DC04 with a width of 50 mm and a thickness of 0.5 mm. The voltage signals from the sensors are digitised using a "CompactRIO" (NI cRIO 9047) with integrated NI 9215 measuring module (analogue voltage input $\pm$ 10 V). Data recording is started via an inductive proximity switch when the press ram passes below a defined stroke height during the stroke movement and is stopped again as it passes the inductive proximity switch during the return stroke movement. Due to the varying process speed caused by the stroke speeds that have been set, the recorded time series have a different number of data points. Further, there are slight variations in the length of the time series withing one experiment. For this reason, all time series are interpolated to a length of 4096 data points and we discard any time series that deviate considerably from the mean length of a run (i.e., threshold of 3). A total of 12 series of experiments, shown in Tab. 9, were carried out with production rates from 80 to 225 spm. To simulate a defect, the spring hardness of the blank holder was manipulated in the test series that were marked as defect. The manipulated experiments result in the component bursting and tearing during production. In a real production environment, this would lead directly to the parts being rejected.
B.3 Data characteristics
@@ B.4 Confounders
In the experiment setting of Tab. 4, the training data set is selected in such a way that the stroke rate correlates with the class label, i.e., there are only normal experiments with small stroke rates and defect ones with high stroke rate. Experiment 1, 2, 3, 10, 11, 12 are the training data and the remaining experiments are the test data. To get a unconfounded setting where the model is not affected by any confounder, we use normal and defect experiments with the same speed in training and respectively test data. This results in experiments 1, 3, 5, 7, 9, 11 in the training set and the remaining in the test set.